A simple pendulum makes 117 complete oscillations in 3.10 min at a location where g = 9.80 m/s2.

(a) Find the period of the pendulum.
1.59 s
(b) Find the length of the pendulum.
m
Answer #1

(a) Total time = 3.1*60 = 186 sec

Time period of the pendulum = 186/117 = 1.59 sec

(b) Time period of pendulum is given by, T = 27

1.59 = 2 V9.80

1 = (1.59/27)* 9.80

l = 0.63 m

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.